# A Reference Architecture for Governance of Cloud Native Applications

### v2

@@ Abstract
The evolution of cloud computing has given rise to Cloud Native Applications (CNAs), presenting new challenges in governance, particularly when faced with strict compliance requirements. This work explores the unique characteristics of CNAs and their impact on governance. We introduce a comprehensive reference architecture designed to streamline governance across CNAs along with a sample implementation, offering insights for both single and multi-cloud environments. Our architecture seamlessly integrates governance within the CNA framework, adhering to a “battery-included” philosophy. Tailored for both expansive and compact CNA deployments across various industries, this design enables cloud practitioners to prioritize product development by alleviating the complexities associated with governance. In addition, it provides a building block for academic exploration of generic CNA frameworks, highlighting their relevance in the evolving cloud computing landscape.
@@ I Introduction
As the adoption of cloud services continues to rise, both cloud Service Providers (CSPs) and Cloud Service Consumers (CSCs) are facing increasingly complex challenges in managing a vast array of resources. This complexity has led to the adoption of advanced observability tools like New Relic [4], Dynatrace [5], and Datadog [6]. However, these tools often fail to provide a holistic view of cloud operations, forcing organizations to rely on multiple solutions which, unfortunately, leads to fragmented data management [7]. Monitoring and observability tools often exhibit weaknesses in addressing compliance and governance requirements, while governance-focused tools tend to lack robust capabilities for monitoring and observability. As a result, both CSPs and CSCs are increasingly inclined to develop their in-house solutions [8]. These solutions, which involve the effective collection, manipulation, storage, and analysis of logs, metrics, and traces, aim to bridge these gaps and provide comprehensive support for monitoring, observability, compliance, and overall cloud governance within cloud environments.
The National Institute of Standards and Technology (NIST) interprets cloud governance as a standardization focus, akin to Internet governance, aiming to ensure interoperability and data portability across various cloud platforms [9]. This perspective anticipates future standards for efficient data and service access tailored to CSCs’ needs [10]. Red Hat describes cloud governance as the framework of policies guiding an organization’s cloud operations [11]. AWS defines cloud governance as a system allowing customers to set and monitor their cloud requirements in terms of security, cost, and oversight [12]. Azure views governance as a tool for safe cloud adoption, balancing control, security, and compliance with cloud scalability [13]. Google cloud specializes its governance definition to data management, ensuring data security, privacy, and compliance with internal and external standards [14]. IBM considers governance as practices to balance benefits, risk, and resources in cloud environments [15, 16]. In this work, governance encompasses the management of all aspects of cloud computing and applications, including CNAs, covering functions like monitoring, observability, and compliance for comprehensive cloud management.
@@ I-B The Rise of CNAs
Inspired by the evolution of cloud computing, a new breed of cloud applications has evolved. Being native to cloud, CNAs are inherently designed, implemented, and maintained to leverage the unique capabilities of cloud technologies.
The Cloud Native Computing Foundation (CNCF) defines CNAs as applications specifically built to utilize the innovations in cloud computing. The CNCF emphasizes that only applications designed with cloud computing’s offerings and features can fully exploit the cloud’s inherent resiliency and scalability [23]. Given their increasing importance and popularity, CNAs have emerged as a critical focus area in the cloud computing. However, the absence of comprehensive governance architectures for CNAs presents significant risks in their design and implementation [24, 25, 26]. This paper addresses this gap by introducing a specialized reference architecture aimed at the effective governance of CNAs. For readers seeking a more in-depth exploration of CNAs and their characteristics, please refer to Section II-A for further details.
@@ II-A3 Serverless as computing platform
CNAs are composed of various components with different run-time requirements. For instance, while the user-interface components and Online Transaction Processing services need to be up and running almost $24/7$, batch processing tasks, aggregations, and backup operations demand a shorter and less frequent run-time. Thus, dedicating a VM or container to a service that runs once or a few times a day is an infeasible under-utilization [41]. The better utilization of resources on the CSPs translates to a more feasible solution for the CSCs [42]. CSPs have brought up a new computing platform known as serverless to address schedule-able workloads. Serverless computing hides server usage from developers and runs the associated code on-demand, only charging the user based on the time that code is running [43]. Interestingly, serverless computing and microservices have a lot of common characteristics. If a microservice does not need to be up and running at all times, a serverless computing platform can bring even more financial feasibility than containers [44, 45]. CSPs have developed and released services such as AWS Lambda [46], Azure Serverless [47], Google cloud Serverless Computing [48], and IBM Cloud Code Engine [49].
@@ II-A4 State isolation
CNAs are built and implemented with clear separation between stateless and stateful services. In the context of CNAs, even if stateless services interact with stateful services, this interaction is often managed via a set of exposed APIs [50]. This separation allows different scalability practices for each stateless and stateful component. Stateless components, often implemented as containers or serverless functions, require less effort for scalability and reliability. In contrast, stateful component scalability and reliability are not as straightforward and require architectural consideration, including replication and storing the state at any given time. An isolated state is one of the desired characteristics of the CNAs [32, 51].
@@ II-B6 Infrastructure governance challenges
Deploying CNAs effectively on Infrastructure as a Service (IaaS) and Platform as a Service (PaaS) demands thorough infrastructure governance, which is central to comprehensive cloud governance. In these environments, practitioners must incorporate governance solutions, key for monitoring and managing CNA deployments. This governance aspect entails synthesizing diverse telemetry data from numerous sources to create an accurate representation of the CNA’s operational state. The complexity of managing CNA infrastructure, given the variety of data sources and the depth of information, is critical. Previous research has explored into these challenges, underlining the evolving demands of infrastructure monitoring in the context of CNAs [62, 66, 65].
@@ III-C Desired Construction, Deployment, and Maintenance Requirements
For effective governance across the lifecycle of CNAs, a comprehensive reference architecture must fulfill specific construction, deployment, and maintenance requirements. These requirements, drawn from extensive field experience and inspired by existing cloud-native architectures [97, 98, 99, 100] ensure that the architecture supports a wide range of governance activities effectively. They address the essential aspects of adaptability, resilience, and efficient management crucial for CNAs operating in dynamic cloud environments.
@@ III-C4 XaC and IaC
In an effective governance framework, XaC and IaC are necessary for managing system components through code, improving traceability and automation across cloud computing. IaC, as part of XaC, plays a key role in CNAs management by facilitating seamless delivery, and deployment and ensuring high levels of automation [105, 106, 107]. These approaches offer key benefits such as consistency, scalability, version control, auditability, and portability, critical for various code-as-policy implementations, including IaC, a crucial feature for CNAs [108, 109]. This framework leverages XaC and IaC principles, utilizing platforms like GitHub for source code management [110] and tools like Terraform for resource orchestration [28] to achieve a comprehensive and automated governance model.
@@ III-D Desired Components for Governance of CNAs
To achieve comprehensive governance of CNAs, it is necessary to identify desired components within a reference architecture that are crucial for maintaining CNAs’ integrity, security, and operational efficiency. Recognizing that some components are inherently composite, combining multiple sub-components to cover various aspects of governance, this discussion provides an in-depth examination of each to clarify their individual and combined contributions. By integrating these desired components, the framework aims to foster a robust and effective governance system, ensuring a resilient and adaptive governance structure tailored to the unique needs of CNAs. The list of desired components are: 1. Data ingestion, 2. Fault tolerant data bus, 3. Data manipulation, 4. Data storage (mutable and immutable), and 5. Data analytics. The details are given below.
@@ III-D2 Fault tolerant data bus
In the governance of CNAs, a Fault tolerant data bus is a desired component, essential for managing large volumes of cloud-generated telemetry and facilitating efficient inter-service communication [66]. This component, incorporating both a primary and an auxiliary data bus, ensures resilience and continuous operation, which is especially crucial for monitoring, observability, and compliance within CNAs. The primary data bus, typically employing Pub/Sub or MQ models, is supported by major CSPs, including AWS Simple Queue Service [121], Azure Service Bus [122], Google cloud Pub/Sub [123], and IBM Cloud Message Queue [124]. In events of service disruptions, the auxiliary data bus offers a robust alternative, ensuring that governance processes remain uninterrupted. Designed to operate seamlessly across VMs, containers, and serverless platforms, this dual data bus system employs peer-to-peer networks for enhanced flexibility and resilience, which is critical for the dynamic environments of CNAs. This setup underscores the importance of fault tolerance in maintaining a consistent and reliable governance framework. Gossip-based peer-to-peer networks [125] or more specialized versions for content distribution [126] can be employed, enhancing flexibility and resilience across different cloud scenarios. This strategic approach ensures that governance processes remain continuous and uninterrupted, even in the face of primary data transfer service disruptions. For simplicity, in the rest of this paper, we will refer to these two types of data buses as Data bus.
@@ III-D5 Data analytics
The diversity in the types of assessments conducted by these components-explicit, range, and baseline-demonstrates their comprehensive approach to data analysis. Explicit assessments ensure compliance with predefined criteria, such as verifying the correct number of nodes in a Kubernetes cluster. Range assessments monitor metrics to stay within established limits, like maintaining acceptable database response times. The most complex baseline assessments develop a model of the ‘desired’ state to identify and address deviations, underlining the necessity for adaptive governance [65, 64, 68]. These analytics components are indispensable for recognizing deviations from governance standards, including SLAs and configurations, by analyzing data to ascertain if a resource aligns with its defined ideal state. Detected discrepancies prompt immediate actions, enabling proactive governance. Our prior research attests to the efficacy of these data analytics in analyzing cloud-generated data in real-time, underscoring their indispensable role in a governance framework. This adaptability and specialization ensure that the architecture can meet the intricate requirements of cloud governance effectively, highlighting the importance of data analytics as a desired component for comprehensive CNA governance.
@@ IV-B Aligning Governance Components with the Proposed Architecture
The second component in our governance architecture is the Fault tolerant data bus. This component is deployed within RGs $1$ to $n$ and in RG-GOV-IMS. The inclusion of these data buses is crucial for ensuring efficient data flow and governance throughout the entire cloud environment. They serve as critical conduits for transmitting telemetry data and governance-related information. While our architecture diagram distinguishes the data buses in RGs and RG-GOV-IMS for clarity, the practical implementation may involve utilizing a singular data bus service, should the CSP offer such a unified solution. For instance, services like AWS Kinesis [146] or IBM Event Streams [147] could be employed to streamline the data flow processes. This unified approach not only simplifies the data management but also enhances the governance framework’s efficiency by reducing the complexity and potential points of failure in the data transmission paths.
@@
The fifth desired component of our reference architecture is the Data analytics. This composite component consists of many data analytics instances, and they are more than just data analysis tools; rather, they encompass specific business logic aimed at addressing diverse aspects of cloud governance, including monitoring, observability, and compliance. Each Data analytics component is designed to be focused and atomic, targeting particular governance needs and integrating smoothly into the overall governance framework. To host these Data analytics components, a dedicated RG, RG-GOV-DA , is established. This allocation highlights their importance in identifying deviations from established governance standards and criteria. The data analytics components are essential for analyzing incoming data, ensuring that resources align with their defined desired states, and triggering alerts for proactive governance actions. This design element strengthens the reference architecture’s capability to manage the complex and diverse demands of cloud governance.
@@ IV-D Architectural Pattern
Finally, we summarize information about our reference architecture using an architectural pattern template [141, Sec. 13.1] as follows.
@@ IV-D2 Problem Description
CNAs’ core characteristics, like heterogeneous computing platforms and microservice architecture, present significant governance challenges. Each computing platform demands specific methods for log collection and analysis. It is crucial to have a detailed view of each component and a holistic view of the entire platform to accurately depict the system’s current state. These challenges include managing the distributed nature of microservices, limitations in serverless platforms, and the dynamic aspects of containers. There is also the need to handle diverse telemetry data efficiently, ensuring clean, formatted, and cross-referenced data to reflect an accurate system state. Moreover, governance in CNAs must account for Big Data characteristics inherent in telemetry, demanding advanced solutions for effective storage and analysis. These complexities, detailed in Section II-B, underscore the need for a sophisticated and comprehensive governance framework that can address the unique hurdles presented by each architectural element and data type within CNAs.
@@ V Real-world Instantiation of the Proposed Reference Architecture
Cloud-native applications (CNAs) are containerized using Docker and deployed on both AWS and IBM. These CNAs emit telemetry data, including metrics and logs, every second for a total of 4,000 entries. AWS employs additional components, such as API Gateways and SQS, to handle the high velocity of data, while IBM utilizes a streamlined approach to transmit telemetry directly to governance pipelines.
@@
We demonstrate the system’s ability to maintain performance consistency across cloud providers. For comprehensive implementation details, including deployment instructions, telemetry formats, data flow, time-delay analytics, and cross-cloud performance metrics, refer to Appendix C. The GitHub repository [161] contains example code, deployment scripts, and documentation structured around GitOps principles, facilitating modular, reproducible multi-cloud deployments.

### v3

@@ Abstract
The evolution of cloud computing has given rise to Cloud Native Applications (CNAs), presenting new challenges in governance, particularly when faced with strict compliance requirements. This work explores the unique characteristics of CNAs and their impact on governance. We introduce a comprehensive reference architecture designed to streamline governance across CNAs, along with a sample implementation, offering insights for both single and multi-cloud environments. Our architecture seamlessly integrates governance within the CNA framework, adhering to a “battery-included” philosophy. Tailored for both expansive and compact CNA deployments across various industries, this design enables cloud practitioners to prioritize product development by alleviating the complexities associated with governance. In addition, it provides a building block for academic exploration of generic CNA frameworks, highlighting their relevance in the evolving cloud computing landscape.
@@ I Introduction
The National Institute of Standards and Technology (NIST) interprets cloud governance as a standardization focus, akin to Internet governance, aiming to ensure interoperability and data portability across various cloud platforms [4]. This perspective anticipates future standards for efficient data and service access tailored to CSCs’ needs [5]. Red Hat describes cloud governance as the framework of policies guiding an organization’s cloud operations [6]. AWS defines cloud governance as a system allowing customers to set and monitor their cloud requirements in terms of security, cost, and oversight [7]. Azure views governance as a tool for safe cloud adoption, balancing control, security, and compliance with cloud scalability [8]. Google Cloud specializes in its governance definition of data management, ensuring data security, privacy, and compliance with internal and external standards [9]. IBM considers governance as practices to balance benefits, risk, and resources in cloud environments [10, 11]. In this work, governance encompasses the management of all aspects of cloud computing and applications, including CNAs, covering functions like monitoring, observability, and compliance for comprehensive cloud management.
@@ I-B The Rise of CNAs
CNAs have emerged as a modern class of cloud applications, inherently designed to leverage the unique capabilities of cloud environments. The Cloud Native Computing Foundation (CNCF) defines CNAs as software systems specifically built to exploit cloud innovations, emphasizing their potential to deliver resiliency and scalability when designed accordingly [18]. Given their increasing importance, CNAs have become a critical focus area in cloud computing. However, the lack of comprehensive governance architectures presents significant risks in their design and implementation [19, 20, 21]. This paper introduces a reference architecture to address these governance challenges. For a more detailed discussion of CNA characteristics, please refer to Section II-A.
@@ I-D Our Contribution
To address these research questions, we employed a methodology grounded in software engineering principles, specifically through the lens of functional and non-functional requirements analysis. Our approach involved systematically analyzing the challenges presented by each RQ and iteratively designing a governance-oriented reference architecture. This methodology enabled us to balance theoretical insights with practical design considerations while ensuring that the resulting architecture remains extensible, implementation-agnostic, and suitable for multi-cloud environments.
@@ II-A3 Serverless as computing platform
CNAs are composed of various components with different run-time requirements. For instance, while the user-interface components and Online Transaction Processing services need to be up and running almost $24/7$, batch processing tasks, aggregations, and backup operations demand a shorter and less frequent run-time. Thus, dedicating a VM or container to a service that runs once or a few times a day is an infeasible under-utilization [36]. The better utilization of resources on the CSPs translates to a more feasible solution for the CSCs [37]. CSPs have brought up a new computing platform known as serverless to address schedulable workloads. Serverless computing hides server usage from developers and runs the associated code on-demand, only charging the user based on the time that code is running [38]. Interestingly, serverless computing and microservices have a lot of common characteristics. If a microservice does not need to be up and running at all times, a serverless computing platform can bring even more financial feasibility than containers [39, 40]. CSPs have developed and released services such as AWS Lambda [41], Azure Serverless [42], Google cloud Serverless Computing [43], and IBM Cloud Code Engine [44].
@@ II-A4 State isolation
CNAs are built and implemented with a clear separation between stateless and stateful services. In the context of CNAs, even if stateless services interact with stateful services, this interaction is often managed via a set of exposed APIs [45]. This separation allows different scalability practices for each stateless and stateful component. Stateless components, often implemented as containers or serverless functions, require less effort for scalability and reliability. In contrast, stateful component scalability and reliability are not as straightforward and require architectural consideration, including replication and storing the state at any given time. An isolated state is one of the desired characteristics of the CNAs [27, 46].
@@ II-B6 Infrastructure governance challenges
Deploying CNAs effectively on Infrastructure as a Service (IaaS) and Platform as a Service (PaaS) demands thorough infrastructure governance, which is central to comprehensive cloud governance. In these environments, practitioners must incorporate governance solutions, key for monitoring and managing CNA deployments. This governance aspect entails synthesizing diverse telemetry data from numerous sources to create an accurate representation of the CNA’s operational state. The complexity of managing CNA infrastructure, given the variety of data sources and the depth of information, is critical. Previous research has explored these challenges, underlining the evolving demands of infrastructure monitoring in the context of CNAs [57, 61, 60].
@@ III-C Desired Construction, Deployment, and Maintenance Requirements
For effective governance across the lifecycle of CNAs, a comprehensive reference architecture must fulfill specific construction, deployment, and maintenance requirements. These requirements, drawn from extensive field experience and inspired by existing cloud-native architectures [92, 93, 94, 95], ensure that the architecture supports a wide range of governance activities effectively. They address the essential aspects of adaptability, resilience, and efficient management crucial for CNAs operating in dynamic cloud environments.
@@ III-C4 XaC and IaC
In an effective governance framework, XaC and IaC are necessary for managing system components through code, improving traceability and automation across cloud computing. IaC, as part of XaC, plays a key role in CNAs management by facilitating seamless delivery and deployment and ensuring high levels of automation [100, 101, 102]. These approaches offer key benefits such as consistency, scalability, version control, auditability, and portability, critical for various code-as-policy implementations, including IaC, a crucial feature for CNAs [103, 104]. This framework leverages XaC and IaC principles, utilizing platforms like GitHub for source code management [105] and tools like Terraform for resource orchestration [23] to achieve a comprehensive and automated governance model.
@@ III-D Desired Components for Governance of CNAs
To achieve comprehensive governance of CNAs, it is necessary to identify desired components within a reference architecture that are crucial for maintaining CNAs’ integrity, security, and operational efficiency. Recognizing that some components are inherently composite, combining multiple sub-components to cover various aspects of governance, this discussion provides an in-depth examination of each to clarify their individual and combined contributions. By integrating these desired components, the framework aims to foster a robust and effective governance system, ensuring a resilient and adaptive governance structure tailored to the unique needs of CNAs. The list of desired components are: 1. Data ingestion, 2. Fault-tolerant data bus, 3. Data manipulation, 4. Data storage (mutable and immutable), and 5. Data analytics. The details are given below.
@@ III-D2 Fault tolerant data bus
In the governance of CNAs, a Fault tolerant data bus is a desired component, essential for managing large volumes of cloud-generated telemetry and facilitating efficient inter-service communication [61]. This component, incorporating both a primary and an auxiliary data bus, ensures resilience and continuous operation, which is especially crucial for monitoring, observability, and compliance within CNAs. The primary data bus, typically employing Pub/Sub or MQ models, is supported by major CSPs, including AWS Simple Queue Service [116], Azure Service Bus [117], Google Cloud Pub/Sub [118], and IBM Cloud Message Queue [119]. In the event of service disruptions, the auxiliary data bus offers a robust alternative, ensuring that governance processes remain uninterrupted. Designed to operate seamlessly across VMs, containers, and serverless platforms, this dual data bus system employs peer-to-peer networks for enhanced flexibility and resilience, which is critical for the dynamic environments of CNAs. This setup underscores the importance of fault tolerance in maintaining a consistent and reliable governance framework. Gossip-based peer-to-peer networks [120] or more specialized versions for content distribution [121] can be employed, enhancing flexibility and resilience across different cloud scenarios. This strategic approach ensures that governance processes remain continuous and uninterrupted, even in the face of primary data transfer service disruptions. For simplicity, in the rest of this paper, we will refer to these two types of data buses as Data bus.
@@ III-D5 Data analytics
The diversity in the types of assessments conducted by these components—explicit, range, and baseline—demonstrates their comprehensive approach to data analysis. Explicit assessments ensure compliance with predefined criteria, such as verifying the correct number of nodes in a Kubernetes cluster. Range assessments monitor metrics to stay within established limits, like maintaining acceptable database response times. The most complex baseline assessments develop a model of the ‘desired’ state to identify and address deviations, underlining the necessity for adaptive governance [60, 59, 63]. These analytics components are indispensable for recognizing deviations from governance standards, including SLAs and configurations, by analyzing data to ascertain if a resource aligns with its defined ideal state. Detected discrepancies prompt immediate actions, enabling proactive governance. Our prior research attests to the efficacy of these data analytics in analyzing cloud-generated data in real-time, underscoring their indispensable role in a governance framework. This adaptability and specialization ensure that the architecture can meet the intricate requirements of cloud governance effectively, highlighting the importance of data analytics as a desired component for comprehensive CNA governance.
@@ IV-B Aligning Governance Components with the Proposed Architecture
The second component in our governance architecture is the Fault-tolerant data bus. This component is deployed within RGs $1$ to $n$ and in RG-GOV-IMS. The inclusion of these data buses is crucial for ensuring efficient data flow and governance throughout the entire cloud environment. They serve as critical conduits for transmitting telemetry data and governance-related information. While our architecture diagram distinguishes the data buses in RGs and RG-GOV-IMS for clarity, the practical implementation may involve utilizing a singular data bus service, should the CSP offer such a unified solution. For instance, services like AWS Kinesis [141] or IBM Event Streams [142] could be employed to streamline the data flow processes. This unified approach not only simplifies the data management but also enhances the governance framework’s efficiency by reducing the complexity and potential points of failure in the data transmission paths.
@@
The fifth desired component of our reference architecture is the Data analytics. This composite component consists of many data analytics instances, and they are more than just data analysis tools; rather, they encompass specific business logic aimed at addressing diverse aspects of cloud governance, including monitoring, observability, and compliance. Each Data analytics component is designed to be focused and atomic, targeting particular governance needs and integrating smoothly into the overall governance framework. To host these Data analytics components, a dedicated RG, RG-GOV-DA is established. This allocation highlights their importance in identifying deviations from established governance standards and criteria. The data analytics components are essential for analyzing incoming data, ensuring that resources align with their defined desired states, and triggering alerts for proactive governance actions. This design element strengthens the reference architecture’s capability to manage the complex and diverse demands of cloud governance.
@@ IV-D Architectural Pattern
Finally, we summarize information about our reference architecture using an architectural pattern template [136, Sec. 13.1] as follows.
@@ IV-D2 Problem Description
CNAs’ core characteristics, like heterogeneous computing platforms and microservice architecture, present significant governance challenges. Each computing platform demands specific methods for log collection and analysis. It is crucial to have a detailed view of each component and a holistic view of the entire platform to accurately depict the system’s current state. These challenges include managing the distributed nature of microservices, limitations in serverless platforms, and the dynamic aspects of containers. There is also a need to handle diverse telemetry data efficiently, ensuring clean, formatted, and cross-referenced data to reflect an accurate system state. Moreover, governance in CNAs must account for Big Data characteristics inherent in telemetry, demanding advanced solutions for effective storage and analysis. These complexities, detailed in Section II-B, underscore the need for a sophisticated and comprehensive governance framework that can address the unique hurdles presented by each architectural element and data type within CNAs.
@@ IV-D3 Solution Description
Additionally, while the modular nature of the architecture allows for flexibility, it may introduce integration complexity, particularly when aligning existing tooling with governance pipelines. In high-throughput CNA environments—where telemetry data may be generated at extreme volumes—certain components such as real-time ingestion or analytics layers may become bottlenecks if not scaled appropriately. Moreover, because the architecture offers a generalized reference model rather than prescriptive implementation steps, adopters must invest effort into aligning it with specific compliance regimes, organizational policies, or performance thresholds. This can result in increased time-to-deployment or additional resourcing for architecture customization and validation.
@@ IV-F Integrating and Advancing Existing Solutions for Enhanced Cloud Governance
Many governance tools provide standardized mechanisms to either push data (e.g., through webhooks or streaming APIs) or allow periodic data pulls (e.g., via scheduled queries or REST APIs). These mechanisms are typically backed by persistent storage layers, including relational databases, NoSQL engines, or other storage systems. The proposed architecture can accommodate these data exchange strategies by aligning them with the ingestion pathways defined in the Data collector agents and Data manipulation stages within the RG-GOV-IMS resource group. These integration points make it feasible to import telemetry, audit trails, or configuration data from a variety of third-party tools, providing both observability and compliance enrichment while avoiding redundant tooling or duplicated pipelines.
@@ V Real-world Instantiation of the Proposed Reference Architecture
CNAs are containerized using Docker and deployed on both AWS and IBM. These CNAs emit telemetry data, including metrics and logs, every second for a total of 4,000 entries. AWS employs additional components, such as API Gateways and SQS, to handle the high velocity of data, while IBM utilizes a streamlined approach to transmit telemetry directly to governance pipelines.
@@
We demonstrate the system’s ability to maintain performance consistency across cloud providers. A graphical representation of the data flow within the instantiated reference architecture is available in Appendix C, providing a visual walkthrough of the end-to-end operational workflow. This deployment also serves as a practical case study to guide cloud and cloud-native adopters in implementing governance architectures tailored to their own environments. Recognizing that no two CNA deployments are alike, the proposed framework is intentionally designed for flexibility, allowing organizations to incorporate additional components or modify existing ones to meet their specific needs.
For comprehensive implementation details, including deployment instructions, telemetry formats, data flow, time-delay analytics, and cross-cloud performance metrics, refer to Appendix C.
The GitHub repository [156] contains example code, deployment scripts, and documentation structured around GitOps principles, facilitating modular, reproducible multi-cloud deployments.